# MoGe: Unlocking Accurate Monocular Geometry Estimation for Open-Domain Images with Optimal Training Supervision

### v1

@@ A.2 Optimal alignment
$\begin{split}0<&l_{1}^{\prime+}(s^{*})-l_{1}^{\prime-}(s^{*})\\ =&2\sum_{{\hat{x}_{i}\over x_{i}}=s^{*}}w_{i}x_{i}-\!\!\!\!\!\!\sum_{{w_{i}\hat{x}_{i}-\tau\over w_{i}x_{i}}=s^{*}}\!\!\!\!w_{i}x_{i}-\!\!\!\!\!\!\sum_{{w_{i}\hat{x}_{i}+\tau\over w_{i}x_{i}}=s^{*}}\!\!\!\!w_{i}x_{i}\\ <&2\sum_{{\hat{x}_{i}\over x_{i}}=s^{*}}w_{i}x_{i}\end{split}$ (17)
Lemma 1 has proved that the minimum point of the objective function in Eq. 13 still occurs among $\left\{x_{i}\over\hat{x}_{i}\right\}$, although the function is non-convex and may contain local minima. Therefore, it takes two steps to solve the subproblem. First, find all extrema among $\left\{x_{i}\over\hat{x}_{i}\right\}$ such that $l_{0}^{\prime-}\left(x_{i}\over\hat{x}_{i}\right)<0\leq l_{0}^{\prime+}\left(x_{i}\over\hat{x}_{i}\right)$ by evaluating the derivative values, which takes $O(N\log N)$ time. Second, evaluate the objective values of those extrema and then find the minimum, which takes $O(Nn_{e})$ time. $n_{e}$ is the number of extrema, approximating the number of outliers , which is expected to be a small constant in practice.
Alignment with 1D shift
Recall the objective in Eq. 4. Let $w_{i}$ be $1/z_{i}$. We apply clipping to each absolute term and rewrite it as follows
$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}\min(\tau,w_{i}|s\hat{x}_{i}\!-x_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|)\end{split}$ (18)
@@
$\begin{split}l_{2}(s,t_{z})=&\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{x}_{i}-x_{i}|}\right)\\ &+\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{y}_{i}-y_{i}|}\right)\\ &+\sum_{i=1}^{n}\min\left(\tau,w_{i}{|t_{z}-(z_{i}-s\hat{z}_{i})|}\right)\end{split}$ (19)
Given arbitrary $s$, using Lemma 1, there exists at least one pair $(t_{z},k)$ such that $t_{z}-(z_{k}-s\hat{z}_{k})=0$ and $t_{z}$ minimizes $\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{z}_{i}+t_{z}-z_{i}|}\right)$, hence minimizes $l(s,t_{z})$ as the rest parts are constant with regard to $t_{z}$. Therefore, a solution $s^{*}$ is always associated with corresponding $(t^{*}_{z},k^{*})$ such that that $s^{*}\hat{z}_{k^{*}}+t^{*}_{z}-z_{k^{*}}=0$. ∎
Lemma 2 allows us to reduce Eq. 18 to the subproblem with respect to some index $k$. For each possible index $k$, the objective is formed as
$\begin{split}\min_{s}\sum_{i=1}^{n}\min(\tau,w_{i}|s\hat{x}_{i}\!-\!sx_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s(\hat{z}_{i}-\hat{z}_{k})-(z_{i}-z_{k})|)\end{split}$ (20)
@@
Alignment with 3D shift
@@
Appendix B Data details
@@ B.1 Training data
Table 6: A table of datasets used for training. #Frames stands for the number of used frames. The types are in short for A. synthetic, B. structure from motion / multiview reconstruction, C. LiDAR / Laser; D. Microsoft Kinect
@@ B.2 Evaluation data
  • NYUv2 [38] We use the official test split of 654 samples Due to the inaccuracy of ground truth values captured by Kinect V1 near boundaries, we filter and remove boundary regions by a simple edge detection method. Specifically, we set a threshold for the difference between the minimum and maximum depth values within a local window. Depth values beyond 5 meters are excluded because they are unreliable due to the limited sensor range. Additionally, we manually mask out areas with reflective and transparent surfaces, such as mirrors and glass, which cannot be accurately captured by the sensor.
@@
  • KITTI [51] We utilizes the test split of 652 images of Eigen et. al. [18] following previous works. The original resolution of $1242\times 375$ does not match our training aspect ratio (ranging from $1:2$ to $2:1$), so we apply a center crop to obtain a $750\times 375$ resolution from the raw images.
@@
  • ETH3D [49] All 454 images are included. The images are undistorted with the official calibration data and downsized from the original $6202\times 4135$ resolution to $2048\times 1365$.
  • iBims-1 [31] All 100 images are included at an original resolution of $640\times 480$.
@@
  • GSO [16] The dataset contains 1,030 objects. For each object, we render a single view at $512\times 512$ resolution. The view is randomly sampled with a FoV ranging from $30^{\circ}$ to $60^{\circ}$. The object is centered in the image, and its bounding box occupies approximately 70% of the image’s size.
@@
  • Sintel [9] We use all 1,064 frames and perform a center crop to $872\times 436$ from the original $1024\times 436$ resolution to fit our aspect ratio range. The sky regions are manually masked out.
@@
  • DDAD [24] We randomly select 1,000 samples from the validation set. The dataset was collected using multiple cameras and LiDAR sensors mounted on a moving vehicle. Some cameras inadvertently capture parts of the vehicle, causing discrepancies with the sensor’s depth data. To address this, we crop the regions that are not obstructed by the vehicle itself.
@@
  • DIODE [52] We utilize the official validation split, which includes 325 indoor images and 446 outdoor images at an original resolution of $1024\times 768$. Due to artifacts in ground truth depth values near the boundaries in this dataset, we identify and remove these boundary regions using a similar approach as described above.
@@ Appendix C More results
C.1 Full table of depth estimation results
In Table 7, we present a full table of results for depth estimation where methods that predict metric or scale-invariant depth are evaluated for affine-invariant depth and disparity for fair comparison with other baselines.
@@ C.1 Full table of depth estimation results
Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.944 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.38
Table 7: Full table of comparison for depth map estimation. * Methods have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results in smaller text size for each version. Gray numbers denote models trained on respective benchmarks.
C.2 Qualitative comparison on uncurated in-the-wild images
We further utilize the first 100 in-the-wild images in DIV2K[2] dataset to extensively compare our method with LeReS [69], UniDepth [41], DUSt3R [57], Metric3D V2 [26] and Depth Anything V2 [66]. The results can be found in the supplementary videos. Since Metric3D V2 and Depth Anything V2 predict depth map and require ground truth camera focal to obtain 3D points cloud results, we visualize them using our estimated focal lengths.
C.3 More visual results
In the following pages, we demonstrate more examples of our results for in-the-wild images. The rows from left to right are the input images, the disparity maps, surface geometry visualization from the source view and three novel view image of the point cloud, respectively.

### v2

@@ A.2 ROE Alignment
$\begin{split}0<&l_{1}^{\prime+}(s^{*})-l_{1}^{\prime-}(s^{*})\\ =&2\sum_{{x_{i}\over\hat{x}_{i}}=s^{*}}w_{i}\hat{x}_{i}-\!\!\!\!\!\!\sum_{{w_{i}x_{i}-\tau\over w_{i}\hat{x}_{i}}=s^{*}}\!\!\!\!w_{i}\hat{x}_{i}-\!\!\!\!\!\!\sum_{{w_{i}x_{i}+\tau\over w_{i}\hat{x}_{i}}=s^{*}}\!\!\!\!w_{i}\hat{x}_{i}\\ \leq&2\sum_{{x_{i}\over\hat{x}_{i}}=s^{*}}w_{i}\hat{x}_{i}.\end{split}$ (17)
Algorithm 3 ROE alignment subproblem w/ truncation
input: arrays $\hat{X}[1..n]$, $X[1..n]$, $W[1..n]$, float $\tau$
output: the optimal scale $s^{*}$, objective value $l^{*}$ to Eq. 13
function SolveSubproblem($\hat{X}$, $X$, $W$, $\tau$)
       $A[1..n]$leftarrow $X/\hat{X}$
       $B[1..n]$leftarrow $(W*X-\tau)/(W*\hat{X})$
       $C[1..n]$leftarrow $(W*X+\tau)/(W*\hat{X})$
       for each array $\mathcal{A}$ in $\{A,B,C\}$ do
           sort $\mathcal{A}$ and obtain sorted indices $I_{\mathcal{A}}[1..n]$
           $Q_{\mathcal{A}}[1..n]$leftarrow accumulated sum of $\{W\hat{X}[I_{\mathcal{A}}[i]]\}_{i=1}^{n}$
       end for
       Initialize $I_{E}$ as empty set
       for $i=1$ to $n$ do $\triangleright$ parallel computation
           for each array $\mathcal{A}$ in $\{A,B,C\}$ do
               $j_{\mathcal{A}}^{-}$leftarrow the last $j$ s.t. $\mathcal{A}[j]<X[i]/\hat{X}[i]$
               $j_{\mathcal{A}}^{+}$leftarrow the last $j$ s.t. $\mathcal{A}[j]\leq X[i]/\hat{X}[i]$
           end for
           $d^{-}$leftarrow $2\cdot Q_{A}[j_{A}^{-}]-Q_{B}[j_{B}^{-}]-Q_{C}[j_{C}^{-}]$
           $d^{+}$leftarrow $2\cdot Q_{A}[j_{A}^{+}]-Q_{B}[j_{B}^{+}]-Q_{C}[j_{C}^{+}]$
           if $d^{-}<0\leq d^{+}$ then append $i$ to $I_{E}$
       end for
       Initialize $l[1..N]$ with $\infty$
       for $i$ in $I_{E}$ do $\triangleright$ parallel computation
           $s$leftarrow $X[i]/\hat{X}[i]$
           $l[i]$leftarrow objective function value at $s$
       end for
       $i^{*}$leftarrow index of the minimum in $l[i],i\in I_{E}$
       $s^{*}$leftarrow $X[i^{*}]/\hat{X}[i^{*}]$, $l^{*}$leftarrow $l[i^{*}]$
       return $s^{*}$, $l^{*}$
end function
Alignment with 1D shift.
Recall the alignment objective and let $w_{i}$ be $1/z_{i}$. We rewrite it as follows:
$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}[w_{i}|s\hat{x}_{i}\!-x_{i}|+w_{i}|s\hat{y}_{i}\!-\!y_{i}|+w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|],\end{split}$ (18)
or apply truncation to each absolute residual term
$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}[\min(\tau,w_{i}|s\hat{x}_{i}\!-x_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|)].\end{split}$ (19)
The proposed solution is outlined in Algorithm 4, with proof as follows. The corresponding subproblem solver is selected based on whether truncation is applied.
@@
$\begin{split}l_{2}(s,t_{z})&=\sum_{i=1}^{n}[\min\left(\tau,w_{i}{|s\hat{x}_{i}-x_{i}|}\right)+\min\left(\tau,w_{i}{|s\hat{y}_{i}-y_{i}|}\right)]\\ &+\sum_{i=1}^{N}\min\left(\tau,w_{i}{|t_{z}-(z_{i}-s\hat{z}_{i})|}\right)].\end{split}$ (20)
Given arbitrary $s$, using Lemma 1, there exists at least one pair $(t_{z},k)$ such that $t_{z}-(z_{k}-s\hat{z}_{k})=0$ and $t_{z}$ minimizes $\sum_{i=1}^{n}\min\left(\tau,w_{i}{|s\hat{z}_{i}+t_{z}-z_{i}|}\right)$, hence minimizes $l_{2}(s,t_{z})$ as the rest parts are constant with regard to $t_{z}$. Therefore, a solution $s^{*}$ is always associated with corresponding $(t^{*}_{z},k^{*})$ such that that $s^{*}\hat{z}_{k^{*}}+t^{*}_{z}-z_{k^{*}}=0$. ∎
Lemma 2 allows us to reduce Eq. 19 to the subproblem with respect to some index $k$. For each possible index $k$, the objective is formed as:
$\begin{split}\min_{s}\sum_{i=1}^{n}\min(\tau,w_{i}|s\hat{x}_{i}\!-\!sx_{i}|)\!+\!\min(\tau,w_{i}|s\hat{y}_{i}\!-\!y_{i}|)\\ +\min(\tau,w_{i}|s(\hat{z}_{i}-\hat{z}_{k})-(z_{i}-z_{k})|),\end{split}$ (21)
@@
Algorithm 4 ROE alignment w/ 1-D shift
input: point arrays $\hat{P}[1..N,1..3]$, $P[1..N,1..3]$,
   weight array $W[1..N]$
output: the optimal scale $s^{*}$, shift $t_{z}^{*}$,
   objective value $l^{*}$ to Eq. 18 or Eq. 19
$W[1..3N]$leftarrow repeat each element in $W$ 3 times
Initialize arrays $s[1..N]$, $l[1..N]$, $t_{z}[1..N]$
for $k=1$ to $N$ do $\triangleright$ parallel computation
[…]
       $(s[k],l[k])$leftarrow SolveSubproblem($\hat{X}$,$X$,$W$)
       $t_{z}[k]$leftarrow $P[k,3]-s[k]\cdot\hat{P}[k,3]$
end for
$k^{*}$leftarrow index of the minimum in $l[1..N]$
$s^{*}$leftarrow $s[k^{*}]$, $l^{*}$leftarrow $l[k^{*}]$, $t_{z}^{*}$leftarrow $t_{z}[k^{*}]$
return $s^{*}$, $t_{z}^{*}$, $l^{*}$
Alignment with 3D shift.
We apply truncation and rewrite the objective as follows:
$\begin{split}\min_{s,t_{z}}\sum_{i=1}^{n}&\min(\tau,w_{i}|s\hat{x}_{i}\!+t_{x}-x_{i}|)\\ &+\min(\tau,w_{i}|s\hat{y}_{i}+t_{y}-y_{i}|)\\ &+\min(\tau,w_{i}|s\hat{z}_{i}+t_{z}-z_{i}|).\end{split}$ (22)
@@
Algorithm 5 ROE alignment w/ 3-D shift
input: point arrays $\hat{P}[1..N,1..3]$, $P[1..N,1..3]$,
   weight array $W[1..N]$
output: the optimal scale $s^{*}$, shift $\mathbf{t}^{*}$,
   objective value $l^{*}$ to Eq. 22
$W[1..3N]$leftarrow repeat each element in $W$ 3 times
Initialize arrays $s[1..N]$, $l[1..N]$, $\mathbf{t}[1..N,3]$
for $k=1$ to $N$ do $\triangleright$ parallel computation
[…]
       $(s[k],l[k])$leftarrow SolveSubproblem($\hat{X}$,$X$,$W$)
       $\mathbf{t}[k]$leftarrow $P[k]-s[k]\cdot\hat{P}[k]$
end for
$k^{*}$leftarrow index of the minimum in $l[1..N]$
$s^{*}$leftarrow $s[k^{*}]$, $l^{*}$leftarrow $l[k^{*}]$, $\mathbf{t}^{*}$leftarrow $\mathbf{t}[k^{*}]$
return $s^{*}$, $\mathbf{t}^{*}$, $l^{*}$
Appendix B Experiment Details
@@ B.1 Training Data
To assign balanced weights to the datasets for training, we compute the retrieval probability of each dataset relative to OpenImagesV7 [32], a large and diverse natural image dataset. Specifically, we leverage DINOv2 [41] to extract feature vectors and calculate the probability that the nearest neighbor of a randomly selected image from OpenImagesV7 is found in each respective training dataset.
@@
Type Label quality Applied losses
A. Synthetic Perfect $\infty$ Dense
B. SfM/MV Recon High $\infty$ Dense&Partial
C. LiDAR/Laser High $\sim 100$m Sparse
D. Kinect Medium $\sim 10$m Dense
Table 5: Datasets used for training (top) and the tailored loss combinations (bottom).
@@ B.2 Evaluation Data
Figure 8: Examples of evaluation data preprocessing: (a) Removing mirror and boundary artifacts from the ground truth depth in NYUv2. (b) Excluding sky regions in Sintel. (c) Removing boundary artifacts from the ground truth depth in DIODE.
@@
  • NYUv2 [39]. We use the official test split of 654 samples. Due to the inaccuracy of ground truth values captured by Kinect V1 near boundaries, we filter and remove boundary regions by a simple edge detection method. Specifically, we set a threshold for the difference between the minimum and maximum depth values within a local window. Depth values beyond 5 meters are excluded because they are unreliable due to the limited sensor range [52]. Additionally, we manually mask out areas with reflective and transparent surfaces, such as mirrors and glass, which cannot be accurately captured by the sensor.
@@
  • KITTI [53]. We utilize the test split of 652 images of Eigen et al. [18] following previous works. The original resolution of $1242\times 375$ does not match our training aspect ratio (ranging from $1:2$ to $2:1$), so we apply center cropping to obtain a resolution of $750\times 375$ from the raw images.
@@
  • ETH3D [50]. All 454 images are included. The images are undistorted with the official calibration data and downsized from the original resolution of $6202\times 4135$ to $2048\times 1365$.
  • iBims-1 [31]. All 100 images are included at an original resolution of $640\times 480$.
@@
  • GSO [16]. The dataset contains 1,030 objects. For each object, we render a single view at $512\times 512$ resolution. The view is randomly sampled with a FOV ranging from $30^{\circ}$ to $60^{\circ}$. The object is centered in the image, and its bounding box occupies approximately 70% of the image’s size.
@@
  • Sintel [9]. We use all 1,064 frames and center-crop the images to $872\times 436$ from the original $1024\times 436$ resolution to fit our aspect ratio range. The sky regions are manually masked out because evaluating models with sky depth included is not meaningful.
@@
  • DDAD [24]. We randomly select 1,000 samples from the validation set. The dataset was collected using multiple cameras and LiDAR sensors mounted on a moving vehicle. Some cameras inadvertently capture parts of the vehicle, causing discrepancies with the sensor’s depth data. To address this issue, we crop the regions that are not obstructed by the vehicle itself.
@@
  • DIODE [54]. We utilize the official validation split, which includes 325 indoor images and 446 outdoor images at an original resolution of $1024\times 768$. Due to artifacts in ground truth depth values near the boundaries in this dataset, we identify and remove these boundary regions using a similar approach as described above.
B.3 Evaluation Protocol
For all our models and baselines, predictions and ground truth are aligned in scale (and shift, if applicable) for each image before measuring errors. To clarify the notations in this section:
  • $\hat{\mathbf{p}}_{i}$ and $\mathbf{p}_{i}$ are the predicted and ground-truth points, respectively.
  • $\hat{z}_{i}$ and $z_{i}$ are the predicted and ground-truth depths, which are the $Z$-coordinate of corresponding points.
  • $\mathcal{M}$ is the mask of valid ground-truth.
  • $a$ and $b$ denote the scale and shift used to align predictions with the ground truth for evaluation, to avoid confusion with similar symbols used in the training objectives.
  • Scale-invariant point map. The scale $a^{*}$ to align prediction with ground truth is computed as:
    $a^{*}=\mathop{\arg\!\min}_{a}\sum_{i\in\mathcal{M}}{1\over z_{i}}\|a\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|_{1},$ (23)
  • Affine-invariant point map. The scale $a^{*}$ and shift $\mathbf{b}^{*}$ are computed as:
    $(a^{*},\mathbf{b}^{*})=\mathop{\arg\!\min}_{a,\mathbf{b}}\sum_{i\in\mathcal{M}}{1\over z_{i}}\|a\hat{\mathbf{p}}_{i}+\mathbf{b}-\mathbf{p}_{i}\|_{1}.$ (24)
  • Scale-invariant depth map. The scale $a^{*}$ is computed as
    $a^{*}=\mathop{\arg\!\min}_{s}\sum_{i\in\mathcal{M}}{1\over z_{i}}|a\hat{z}_{i}-z_{i}|.$ (25)
  • Affine-invariant depth map. The scale $a^{*}$ and shift $b^{*}$ are computed as
    $(a^{*},b^{*})=\mathop{\arg\!\min}_{s}\sum_{i\in\mathcal{M}}{1\over z_{i}}|a\hat{z}_{i}+b-z_{i}|.$ (26)
  • Affine-invariant disparity map. We follow the established protocol for affine disparity alignment [43], using least-squares to align predictions in disparity space:
    $(a^{*},b^{*})=\mathop{\arg\!\min}_{s}\sum_{i\in\mathcal{M}}(a\hat{d}_{i}+b-d_{i})^{2},$ (27)
    where $\hat{d}_{i}$ is the predicted disparity and $d_{i}$ is the ground truth, defined as $d_{i}=1/z_{i}$. To prevent aligned disparities from taking excessively small or negative values, the aligned disparity is truncated by the inverted maximum depth $1/z_{\text{max}}$ before inversion. The final aligned depth $\hat{z}_{i}^{*}$ is computed as:
    $\hat{z}_{i}^{*}:={1\over\max(a^{*}\hat{d}_{i}+b^{*},1/z_{\text{max}})}.$ (28)
@@ Appendix C More Results
Full table of depth estimation results.
In Table 6, we present detailed results for depth estimation where methods that predict metric or scale-invariant depth are also evaluated on affine-invariant depth and disparity for a fair comparison.
@@
Ours 3.41 98.6 4.08 98.1 3.17 98.8 3.20 98.1 0.94 100 18.4 79.5 8.91 91.5 4.00 97.3 5.76 95.2 1.38
Table 6: Full table of comparison for depth map estimation. *: methods that have multiple model versions available for respective benchmarks, among which the best for each benchmark is chosen for ranking, followed by the detailed results for each version (smaller font size). Gray numbers denote models trained on respective benchmarks.
More qualitative comparisons.
Fig. 10 and Fig. 11 present additional visual comparisons on zero-shot evaluation datasets and in-the-wild images. Our method is compared with LeReS [71], UniDepth [42], DUSt3R [59], Metric3D V2 [26] and Depth Anything V2 [68]. Since Metric3D V2 and Depth Anything V2 predict depth map and require ground truth camera focal to obtain 3D points cloud results, we visualize them using our estimated focal lengths.
In the supplementary videos, we present extensive and uncurated comparisons using the first 100 images from the DIV2K[2] dataset.
More visual results
In Fig. 12 and Fig. 13, we demonstrate more reconstruction results of our method for more open-domain images.
Appendix D Limitations and Future Work
While our model demonstrates strong performance, accurately capturing thin structures remains a significant challenge. This difficulty arises from the network’s limited capacity and the presence of noisy real-world training data. As illustrated in Fig. 9, our model may fail to recover these intricate structures.
Additionally, while monocular video reconstruction holds great promise as an application, achieving temporal consistency presents substantial challenges. Our model, designed for single-image input, cannot inherently maintain temporal coherence due to the ambiguity of the task. Addressing this issue would require non-trivial solutions, such as global optimization techniques. Given the rapid advancements in video depth estimation, we believe that an end-to-end model for monocular video reconstruction could significantly benefit from our proposed techniques. Exploring this direction is a compelling avenue for future work.
Figure 9: A failure case. Our model does not capture the thin structures of the fence, leading to a flattened geometry.
Figure 10: Additional qualitative comparisons from the evaluation datasets. *: for methods without camera intrinsics prediction, ground-truth camera intrinsics (and disparity shifts) were used to lift their results into 3D points. Best viewed with zoom.
Figure 11: Additional qualitative comparisons for in-the-wild images from the DIV2K dataset. *: for methods without camera intrinsics prediction, our camera intrinsics prediction were used to lift their results into 3D points. Supplementary videos contain more results. Best viewed with zoom.
Figure 12: Additional visual results for open-domain images of our model (page 1 of 2). The columns from left to right are the input images, reconstructed disparity maps, reconstructed surface geometry viewed from the source view, and three novel-view images, respectively.
Figure 13: Additional visual results for open-domain images of our model (page 2 of 2). The columns from left to right are the input images, reconstructed disparity maps, reconstructed surface geometry viewed from the source view, and three novel-view images, respectively.